Ali is two years younger than his brother, Rahim. The total of their ages equals 56. How old is each brother?
Answer by macston(5194)

A=Ali; R=Rahim=A+2yrs
A+R=56 yrs Substitute for R.
A+A+2yrs=56 yrs Subtract 2 yrs from each side.
2A=54 yrs Divide each side by 2.
A=27 yrs ANSWER 1: Ali is 27 years old.
R=A+2yrs=27 yrs+2yrs=29 yrs ANSWER 2: Rahim is 29 years old.